1. Madison Corporation sells three products (M, N, and O) in the following mix: 3:1:2. Unit price and cost data are: M N O Unit sales price $ 14 $ 10 $ 13 Unit variable costs 7 6 10 Total fixed costs are $449,500. The selling price per composite unit for the current sales mix (rounded to the nearest cent) is

: Multiple Choice $37.00. $ 12.33. $41.00. $78.00. $52.00.

2.

Shore Company reports the following information regarding its production cost.

Units produced 38,000 units
Direct labor $ 33 per unit
Direct materials $ 34 per unit
Variable overhead $ 290,000 in total
Fixed overhead $ 104,920 in total


Compute product cost per unit under absorption costing.

Multiple Choice

  • $75.00

  • $77.39

  • $67.00

  • $33.00

  • $34.00

3.

A July sales forecast projects that 5,000 units are going to be sold at a price of $12.50 per unit. The management forecasts 2% growth in sales each month. Total August sales are anticipated to be:

Multiple Choice

  • $62,500.

  • $61,250.

  • $65,000.

  • $63,750.

  • $60,000.

4.

Trago Company manufactures a single product and has a JIT policy that ending inventory must equal 20% of the next month's sales. It estimates that May's ending inventory will consist of 56,600 units. June and July sales are estimated to be 283,000 and 293,000 units, respectively. Compute the number of units to be produced that would appear on the company's production budget for the month of June.

Multiple Choice

  • 341,600 units.

  • 283,000 units.

  • 285,000 units.

  • 281,680 units.

  • 226,400 units.

5.

A company uses activity-based costing to determine the costs of its three products: A, B, and C. The budgeted cost and activity for each of the company's three activity cost pools are shown in the following table:

Budgeted Activity
Activity Cost Pool Budgeted Cost Product A Product B Product C
Activity 1 $ 80,000 7,000 10,000 21,000
Activity 2 $ 55,000 8,000 16,000 9,000
Activity 3 $ 102,000 3,500 2,000 2,625


How much overhead will be assigned to Product B using activity-based costing?

Multiple Choice

  • $72,826.99

  • $92,164.37

  • $70,605.13

  • $237,000.00

  • $80,000.00
Answer #1

1.composite units are set of different products that are grouped together in proportion of sales mix, this helps in finding break even units for a company as a whole.

one sale mix is one composite unit.

sales mix m:n:o = 3:1:2

sellprice =M =$14, N =$10, O=$13

=[3*14 + 1*10 + 2*13]

=[42+10+26]

=$78

2.ABSORPTION COSTING CONSIDERS VARIABLE COSTS AS WELL AS FIXED COSTS A PRODUCT COST.

Product cost

Direct material $33
Direct labor $34
Variable overhead [290,000/38000] $7.63
Fixed overhead [104920/38000] $2.76
Total Cost [33+34+7.63+2.76] $77.39

3.July sales =$12.5*5000

=62,500$

August sales = July sales + [july sales*0.02]    (2% growth)

=62500 + [62500*0.02]

=62500+1250

=63750$

4.

June July
Sales 283,000 293,000
Add: desired ending inventory 58600[293000*20%]
Less: Beginning inventory 56600
Production 285000 units[283000+58600-56600]

285000units

5. Predetermined over head = Cost/ activtity

Budgeted Activity
Activity Cost Pool Budgeted Cost Product A Product B Product C total activity
Activity 1 $ 80,000 7,000 10,000 21,000 38000[7000+10000+21000]
Activity 2 $ 55,000 8,000 16,000 9,000 33000[8000+16000+9000]
Activity 3 $ 102,000 3,500 2,000 2,625 8125[3500+2000+2625]

Now will allocated predetermined overhead to Product B

Activity 1 = [80000/38000]*10000=21052.63

Activity 2 =[55000/33000]*16000=26666.66

Activity 3 = [102000/8125]*2000=25107.69

Total = [21052.63+26666.66+25107.69]

72826.99$